OpenStudy (anonymous):

I need help on this problem that has to do with multiplication with elimination. I WILL FAN AND MEDAL YOU!!! 6x - 3y = 12 4x - 3y = 24

OpenStudy (anonymous):

Well you need to get rid of one of the two variables to start so we can subtract to do that 6x - 3y = 12 - 4x - 3y = 24 which ends up being 2x = -12 so x = -6 Plugging this into the equation (second one) 4 * -6 - 3y = 24 -24 - 3y = 24 -3y = 48 y = -16 So the prove, 6 * -6 - 3 * -16 = 12 -36 - -48 = 12 (48 - 36 = 12) 12 = 12 And that's that, or at least it should be.